40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Thu Jan 31 14:00:00 2013, Kristen Holtz, Notes, PMC, PMC Longitudinal Testing and Modal Analysis PMCphoto_circclamps.JPGPMC_ringdown_Jan28.JPGPMC_unconstrainedringdown_Jan28.JPGPMC_Excitation_Diagram_1-28-13.JPG
    Reply  Wed Mar 27 01:52:34 2013, tara, Notes, PMC, PMC Longitudinal Testing and Modal Analysis pmc_eigen.png
       Reply  Sun Mar 31 03:03:28 2013, tara, Notes, PMC, PMC Longitudinal Testing and Modal Analysis psl_log.pngpsl_log2.png
          Reply  Sun Mar 31 14:42:05 2013, Evan, Notes, PMC, PMC Longitudinal Testing and Modal Analysis 
             Reply  Sun Mar 31 20:06:16 2013, tara, Notes, PMC, PMC Longitudinal Testing and Modal Analysis 
       Reply  Mon Apr 1 20:59:19 2013, tara, Notes, PMC, PMC Longitudinal Testing and Modal Analysis 
          Reply  Thu Apr 4 23:55:12 2013, Evan, Notes, PMC, PMC Longitudinal Testing and Modal Analysis mirrorsag_coarse.pdfmirrorsag_fine.pdf
    Reply  Thu Apr 4 11:44:14 2013, Evan, Notes, PMC, 270 Hz clamped PMC twisting mode pmc_clamp_twist_mode.jpg
Message ID: 1097     Entry time: Thu Jan 31 14:00:00 2013     Reply to this: 1131   1138
Author: Kristen Holtz 
Type: Notes 
Category: PMC 
Subject: PMC Longitudinal Testing and Modal Analysis 

Longitudinal Mode frequency @ 420 Hz

 

Monday, January 28, 2013

Tests of resonant frequency specifically for longitudinal mode done on PMC because the piezo will only affect this length.  Originally the PMC spacer was constrained in motion by clamps on the PMC base, as shown below.

PMCphoto_circclamps.JPG

 Data was taken with an oscilloscope and laser vibrometer - the B&K system was not functioning correctly - and the PMC spacer was excited by hitting it with a marker.  The PMC was hit near the endcap, along the long axis of the PMC, as shown in the below image.

PMC_Excitation_Diagram_1-28-13.JPG

 

The results in the time domain are shown below and indicate a resonant frequency of 450 Hz.  A next step is to fit this data to a decaying sine wave.

PMC_ringdown_Jan28.JPG

When the above clamps were loosened to allow for more motion, the frequency dropped to closer to 350 Hz.

PMC_unconstrainedringdown_Jan28.JPG

 

ELOG V3.1.3-